# Nonsmooth Implicit Differentiation: Deterministic and Stochastic Convergence Rates

### v1

@@ Abstract
We study the problem of efficiently computing the derivative of the fixed-point of a parametric non-differentiable contraction map. This problem has wide applications in machine learning, including hyperparameter optimization, meta-learning and data poisoning attacks. We analyze two popular approaches: iterative differentiation (ITD) and approximate implicit differentiation (AID). A key challenge behind the nonsmooth setting is that the chain rule does not hold anymore. Building upon the recent work by Bolte et al. (2022), who proved the linear convergence of non-differentiable ITD, we provide refined linear convergence rates for both ITD and AID in the deterministic case. We further introduce NSID, a new method to compute the implicit derivative when the fixed point is defined as the composition of an outer map and an inner map which is accessible only through a stochastic unbiased estimator. We establish rates for the convergence of NSID to the true derivative, encompassing the best available rates in the smooth setting. We present illustrative experiments confirming our analysis.
@@ 1 Introduction
$T(w,\lambda)$ $=\mathbb{E}[\hat{T}(w(\lambda),\lambda,\xi)],$
@@
where the external map $G$ can be evaluated, but the inner map $T$ is accessible only via a stochastic estimator $\hat{T}$, with $\xi$ a random variable.
@@
A main motivation for computing the implicit derivative of (1) is provided by bilevel optimization, which aims to minimize an upper level objective function of $w(\lambda)$. Important examples are given by hyperparameter optimization and meta-learning (Franceschi et al., 2018; Lee et al., 2019), where (1) expresses the optimality conditions of a lower-level minimization problem. Further examples include learning a surrogate model for data poisoning attacks (Xiao et al., 2015; Muñoz-González et al., 2017), deep equilibrium models (Bai et al., 2019) or OptNet (Amos & Kolter, 2017). All these problems may present nonsmooth mappings $\Phi$. For instance, consider hyperparameter optimization or data poisoning attacks for SVMs, or meta-learning for image classification, where $\Phi$ is evaluated through the forward pass of a neural net with RELU activations (Bertinetto et al., 2019; Lee et al., 2019; Rajeswaran et al., 2019). In addition, when such settings are applied to large datasets, evaluating the map $\Phi$ would be too costly, but we can usually apply stochastic methods through the composite stochastic structure in (2), where only $T$ involves a computation on the full training set (e.g. a gradient descent step).
@@
Nowadays, automatic differentiation techniques (Griewank & Walther, 2008) popular for deep learning, can also be used to efficiently, i.e. with a cost of the same order of that of approximating $w(\lambda)$, approximate Jacobian-vector (or vector-Jacobian) products of $w(\lambda)$ by relying only on an implementation of an iterative solver for problem (1). There are two main approaches to achieve this: ITerative Differentiation (ITD) (e.g. Maclaurin et al. (2015); Franceschi et al. (2017)), which differentiates through the steps of the solver for (1), and Approximate Implicit Differentiation (AID) (e.g. Pedregosa (2016); Lorraine et al. (2020)), which relies on approximately solving the linear system emerging from the implicit expression for the Jacobian-vector product. Despite the analysis of such methods has been usually done in the case that $\Phi$ is smooth, there are now several open source implementations relying on popular deep learning frameworks (e.g. Grazzi et al. (2020); Blondel et al. (2022); Liu & Liu (2021)), which practitioners can use even when $\Phi$ is not differentiable. However, when $\Phi$ is not differentiable despite existing algorithmic proposals (Ochs et al., 2015; Frecon et al., 2018), establishing theoretical convergence guarantees is challenging, since even if the solution map $w$ is almost everywhere differentiable and the Clarke subgradient is well defined, the chain rule of differentiation, exploited by AID and ITD approaches, does not hold.
@@
Recently Bolte & Pauwels (2021) introduced the notion of conservative derivatives as an effective tool to rigorously address automatic differentiation of neural networks with non-differentiable activations (e.g. ReLU). Moreover, if $\Phi(\cdot,\lambda)$ is a contraction and under the general assumption that $\Phi$ if piecewise Lipschitz smooth with finite pieces, Bolte et al. (2022) provide an asymptotic linear convergence rate for deterministic ITD. [^1] Therein, referred to as piggyback automatic differentiation. However, we are not aware of any result of such type for the AID method and for the stochastic setting of problem (2), even when $G(v,\lambda)=v$. In particular the compositional structure (2) allows us to cover e.g. proximal stochastic gradient methods, which are a common and practical example of nonsmooth optimization algorithms, but it adds additional challenges since we do not have access to an unbiased estimator of $\Phi$ as for the smooth stochastic case studied in (Grazzi et al., 2021, 2023).
@@
We present theoretical guarantees on AID and ITD for the approximation of the conservative derivative of the fixed point solution of (1), building upon the analysis of Bolte et al. (2022). Specifically:
  • We prove non-asymptotic linear convergence rates for ITD and AID approaches which extends the results given in (Grazzi et al., 2020) for the case where $\Phi$ is Lipschitz smooth. The given bounds indicate that AID converges faster than ITD, which we verify empirically. We also identify cases in which this difference in performance in favor of AID might be large due to non-differentiable regions.
  • We propose the first stochastic AID approach with proven convergence rates, which we name nonsmooth stochastic implicit differentiation (NSID). Notably, we prove that NSID can converge to the true conservative Jacobian-vector product with rate $O(1/k)$, where $k$ is the number of samples, provided that the fixed-point problem is solved with rate $O(1/k)$.
Finally, we provide experiments on two bilevel optimization problems, hyperparameter optimization and adversarial poisoning attacks, confirming our theoretical finding.
@@
When $\Phi$ is differentiable and under some regularity assumptions, approximation guarantees have been established for AID and ITD approaches in the deterministic setting (Pedregosa, 2016; Grazzi et al., 2020), and for AID approaches in the special case of the stochastic setting (2) where $G(v,\lambda)=v$ and we have access only to $\hat{T}$ (Grazzi et al., 2021, 2023). Furthermore, several works established convergence rates and, in the stochastic setting, sample complexity results for bilevel optimization algorithms relying on AID and ITD approaches, see e.g. (Ghadimi & Wang, 2018; Ji et al., 2021; Arbel & Mairal, 2021; Chen et al., 2021).
Aside from (Bolte et al., 2022), in the nonsmooth case, Bertrand et al. (2020, 2022) present deterministic and sparsity-aware nonsmooth ITD and AID procedures together with asymptotic linear convergence guarantees when $w(\lambda)$ is the solution of a composite minimization problem where one component has a sum structure. Contrary to this work and to (Bolte et al., 2022), their results rely on some differentiability assumptions on the algorithms, which is verified after a finite number of iterations. For bilevel optimization, some recent works have provided stochastic algorithms with convergence rates for the special case where the lower-level problem has linear (Khanduri et al., 2023) or equality (Xiao et al., 2023) constraints.
@@ 2 Preliminaries
$\mathcal{A}\mathcal{C}$ ::=:= $\{AB\,|\,A\in\mathcal{A},C\in\mathcal{C}\}$
$\mathcal{A}^{x}$ ::=:= $\{A^{x}\,|\,A\in\mathcal{A}\}\ \quad\text{with}\ x\in\{\top,-1\}.$
We let $\operatorname{co}(\mathcal{A})$ be the convex envelope of $\mathcal{A}$, and define $\lVert\mathcal{A}\rVert_{\sup}=\sup\{\lVert A\rVert\,|\,A\in\mathcal{A}\}$. For any integer $m\geq 1$ we set $[m]=\{1,\dots,m\}$. If $F\colon\mathbb{R}^{p_{1}+p_{2}}\to\mathbb{R}^{d}$ is a differentiable mapping, we denote by $F^{\prime}(x)\in\mathbb{R}^{d\times(p_{1}+p_{2})}$ the derivative of $F$ (its Jacobian) and by $\partial_{1}F(x)\in\mathbb{R}^{d\times p_{1}}$ and $\partial_{2}F(x)\in\mathbb{R}^{d\times p_{2}}$ the partial derivatives of $F$ with respect to the first and second block of variables respectively. For a random vector $\xi\in\mathbb{R}^{d}$, we denote with $\mathbb{E}[\xi]$ its expectation and with $\operatorname{Var}[\xi]=\mathbb{E}\lVert\xi-\mathbb{E}[\xi]\rVert^{2}$ its variance. In our assumptions we will consider the class of so called definable functions, which includes the large majority of functions used for machine learning applications (see Appendix A).
@@ 2.1 Conservative Derivatives
Given a locally Lipschitz continuous mapping $F\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ defined on the open set $U$, we say that a set-valued mapping $D_{F}\colon U\rightrightarrows\mathbb{R}^{d\times p}$ is a conservative derivative (CD) of $F$, if $D_{F}$ has closed graph, nonempty compact values, and for every absolutely continuous curve $\gamma\colon[0,1]\to\mathbb{R}^{p}$ we have, for almost all $t\in[0,1]$, that
@@
A function $F\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ is called path differentiable if it is locally Lipschtz continuous and has a conservative derivative.
Conservative derivatives are extensively analyzed in (Bolte & Pauwels, 2021). Some key properties are that: (1) they are almost everywhere single-valued and equal to classical derivatives; (2) for path differentiable functions, the Clarke subgradient is the minimal conservative derivative; (3) chain rule holds for conservative derivatives; (4) locally Lipschitz and definable mappings admit conservative derivatives. We also point out that conservative derivatives are not unique: two valid CDs for $F$ can be different in sets of zero Lebesgue measure. This accounts for the fact there are in general multiple ways to express a path differentiable function as a composition of others but applying the chain rule produces always valid CDs.
@@
Note that $\lVert\mathcal{A}\rVert_{\sup}=\operatorname{\it{e}}(\mathcal{A},\{0\})$. The excess satisfies several properties similar to the ones of a distance, even though it is not symmetric (see Lemma B.1). Similarly to (Scholtes, 2012) we give the following definition (which is slightly more general than that given in (Bolte & Pauwels, 2021)).
@@
Let $F_{1},\dots,F_{m}\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ be continuous mappings defined on a nonempty open set $U$. A continuous selection of $F_{1},\dots,F_{m}$ is a continuous mapping $F\colon U\to\mathbb{R}^{d}$ such that for every $x\in\mathbb{R}^{p}\colon F(x)\in\{F_{1}(x),\dots,F_{m}(x)\}$. In such case the active index set mapping is the set-valued mapping $I_{F}\colon U\rightrightarrows[m]$, with $I_{F}(x)=\{i\in[m]\,|\,F_{i}(x)=F(x)\}$. Moreover, if the $F_{i}$’s are differentiable we set $D_{F}^{s}\colon U\subset\mathbb{R}^{p}\rightrightarrows\mathbb{R}^{d\times p}$ such that
@@
Let $F\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ be a continuous selection of definable and continuously differentiable mappings $F_{1},\dots,F_{m}\colon U\to\mathbb{R}^{d}$. Then $F$ is definable if and only if $I_{F}\colon\mathbb{R}^{p}\rightrightarrows[m]$ is definable, and in such case $D_{F}^{s}$ is a conservative derivative of $F$.
We can also define partial conservative derivatives. If $p=p_{1}+p_{2}$ and $F\colon U\subset\mathbb{R}^{p_{1}+p_{2}}\to\mathbb{R}^{d}$, we have $D_{F}\colon U\rightrightarrows\mathbb{R}^{d\times(p_{1}+p_{2})}$ and we set $D_{F,1}\colon U\rightrightarrows\mathbb{R}^{d\times p_{1}}$ and $D_{F,2}\colon U\rightrightarrows\mathbb{R}^{d\times p_{2}}$ such that
$D_{F,1}(x)=\big{\{}A\,|\,[A,B]\in D_{F}^{s}(x)\big{\}}\quad\text{and}\quad D_{F,2}(x)=\big{\{}B\,|\,[A,B]\in D_{F}^{s}(x)\big{\}}.$
Finally, we denote with $F^{\prime}(x)$ an arbitrary element of the conservative derivative $D_{F}(x)$ and by $\partial_{k}F(x)$ an arbitrary element of $D_{F,k}(x)$, which correspond to the classical (partial) derivatives if $F$ is differentiable. By building on (Bolte et al., 2022, Lemma 3), we prove the following result (the proof is in Appendix B).
@@
Let $F\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ be a continuous definable selection of the definable Lipschitz smooth mappings $F_{1},\dots,F_{m}\colon U\to\mathbb{R}^{d}$. Let $L_{i}$ be the Lipschitz constant of $F^{\prime}_{i}$ and set $L=\max_{1\leq i\leq m}L_{i}$. Then for every $x\in\mathbb{R}^{p}$, there exist $R_{x}>0$ such that
$\forall\,x^{\prime}\in U\colon\ \operatorname{\it{e}}(D_{F}^{s}(x^{\prime}),D_{F}^{s}(x))\leq L_{x}(x^{\prime})\lVert x-x^{\prime}\rVert,$ (5)
@@
[…]
Note that in the smooth case ($m=1$), (5) corresponds to global $L$-smoothness (since $M_{x}=0$), while in general it is weaker. In particular, the quantity $L+\frac{M_{x}}{R_{x}}$ is well defined even when $F$ is not differentiable at $x$, but blows up when $x$ approaches a point of non-differentiability, e.g. for $\operatorname{ReLU}(x)=\max(0,x)$, $\lim_{x\to 0^{+}}M_{x}/R_{x}=\infty$.

### v2

@@ Abstract
We study the problem of efficiently computing the derivative of the fixed-point of a parametric nondifferentiable contraction map. This problem has wide applications in machine learning, including hyperparameter optimization, meta-learning and data poisoning attacks. We analyze two popular approaches: iterative differentiation (ITD) and approximate implicit differentiation (AID). A key challenge behind the nonsmooth setting is that the chain rule does not hold anymore. Building upon the recent work by Bolte et al. (2022), who proved linear convergence of nondifferentiable ITD, we provide an improved linear rate for ITD and a slightly better rate for AID, both in the deterministic case. We further introduce NSID, a new stochastic method to compute the implicit derivative when the fixed point is defined as the composition of an outer map and an inner map which is accessible only through a stochastic unbiased estimator. We establish rates for the convergence of NSID, encompassing the best available rates in the smooth setting. We present illustrative experiments confirming our analysis.
@@ 1 Introduction
$T(w,\lambda)$ $=\mathbb{E}[\hat{T}_{\xi}(w(\lambda),\lambda)],$
@@
where the external map $G$ can be evaluated, but the inner map $T$ is accessible only via a stochastic estimator $\hat{T}_{\xi}$, with $\xi$ a random variable.
@@
A main motivation for computing the implicit derivative of (1) is provided by bilevel optimization, which aims to minimize an upper level objective function of $w(\lambda)$. Important examples are given by hyperparameter optimization and meta-learning (Franceschi et al., 2018; Lee et al., 2019), where (1) expresses the optimality conditions of a lower-level minimization problem. Further examples include learning a surrogate model for data poisoning attacks (Xiao et al., 2015; Muñoz-González et al., 2017), deep equilibrium models (Bai et al., 2019) or OptNet (Amos & Kolter, 2017). All these problems may present nonsmooth mappings $\Phi$. For instance, consider hyperparameter optimization or data poisoning attacks for SVMs, or meta-learning for image classification, where $\Phi$ is evaluated through the forward pass of a neural net with RELU activations (Bertinetto et al., 2019; Lee et al., 2019; Rajeswaran et al., 2019). In addition, when such settings are applied to large datasets, evaluating the map $\Phi$ would be too costly, but we can usually apply stochastic methods through the composite stochastic structure in (2), where only $T$ involves a computation on the full training set (e.g., a gradient descent step).
@@
Nowadays, automatic differentiation techniques (Griewank & Walther, 2008) popular for deep learning, can also be used to efficiently, i.e. with a cost of the same order of that of approximating $w(\lambda)$, approximate Jacobian-vector (or vector-Jacobian) products of $w(\lambda)$ by relying only on an implementation of an iterative solver for problem (1). There are two main approaches to achieve this: ITerative Differentiation (ITD) (e.g., Maclaurin et al. (2015); Franceschi et al. (2017)), which differentiates through the steps of the solver for (1), and Approximate Implicit Differentiation (AID) (e.g., Pedregosa (2016); Lorraine et al. (2020)), which relies on approximately solving the linear system emerging from the implicit expression for the Jacobian-vector product. Despite the analysis of such methods has been usually done in the case that $\Phi$ is smooth, there are now several open source implementations relying on popular deep learning frameworks (e.g., Grazzi et al. (2020); Blondel et al. (2022); Liu & Liu (2021)), which practitioners can use even when $\Phi$ is not differentiable. However, when $\Phi$ is not differentiable despite existing algorithmic proposals (Ochs et al., 2015; Frecon et al., 2018), establishing theoretical convergence guarantees is challenging, since even if the solution map $w$ is almost everywhere differentiable and the Clarke subgradient is well defined, the chain rule of differentiation, exploited by AID and ITD approaches, does not hold.
@@
Recently Bolte & Pauwels (2021) introduced the notion of conservative derivatives as an effective tool to rigorously address automatic differentiation of neural networks with nondifferentiable activations (e.g., ReLU). Moreover, if $\Phi(\cdot,\lambda)$ is a contraction and under the general assumption that $\Phi$ is piecewise Lipschitz smooth with finite pieces, Bolte et al. (2022) provide an asymptotic linear convergence rate for deterministic ITD. [^1] Therein, referred to as piggyback automatic differentiation. However, we are not aware of any result of such type for the AID method and for the stochastic setting of problem (2), even when $G(v,\lambda)=v$. In particular the compositional structure (2) allows us to cover e.g., proximal stochastic gradient methods, which are a common and practical example of nonsmooth optimization algorithms, but it adds additional challenges since we do not have access to an unbiased estimator of $\Phi$ as for the smooth stochastic case studied in (Grazzi et al., 2021, 2023).
@@
We present theoretical guarantees on AID and ITD for the approximation of the conservative derivative of the fixed point solution of (1), building upon the framework of Bolte et al. (2022). Specifically:
  • We prove non-asymptotic linear convergence rates for deterministic ITD and AID which, from one hand extend the results for the case where $\Phi$ is Lipschitz smooth given in (Grazzi et al., 2020), and on the other end, improve the result in (Bolte et al., 2022) for nonsmooth ITD. The given bounds indicate that AID converges faster than ITD, which we verify empirically. We also identify cases in which this difference in performance in favor of AID might be large due to nondifferentiable regions.
  • We propose the first stochastic AID approach with proven convergence rates, which we name nonsmooth stochastic implicit differentiation (NSID). Notably, we prove that NSID can converge to a true conservative Jacobian-vector product with rate $O(1/k)$, where $k$ is the number of samples, provided that the fixed-point problem is solved with rate $O(1/k)$.
Finally, we provide experiments on two bilevel optimization problems, i.e. hyperparameter optimization and adversarial poisoning attacks, confirming our theoretical findings.
@@
When $\Phi$ is differentiable and under some regularity assumptions, approximation guarantees have been established for AID and ITD approaches in the deterministic setting (Pedregosa, 2016; Grazzi et al., 2020), and for AID approaches in the special case of the stochastic setting (2) where $G(v,\lambda)=v$ and we have access only to $\hat{T}$ (Grazzi et al., 2021, 2023). Furthermore, several works established convergence rates and, in the stochastic setting, sample complexity results for bilevel optimization algorithms relying on AID and ITD approaches, see e.g., (Ghadimi & Wang, 2018; Ji et al., 2021; Arbel & Mairal, 2021; Chen et al., 2021).
Aside from (Bolte et al., 2022), in the nonsmooth case, Bertrand et al. (2020, 2022) present deterministic and sparsity-aware nonsmooth ITD and AID procedures together with asymptotic linear convergence guarantees when $w(\lambda)$ is the solution of a composite minimization problem where one component has a sum structure. Contrary to this work and to (Bolte et al., 2022), their results rely on some differentiability assumptions on the algorithms, which are verified after a finite number of iterations. For bilevel optimization, some recent works have provided stochastic algorithms with convergence rates for the special case where the lower-level problem has linear (Khanduri et al., 2023) or equality (Xiao et al., 2023) constraints.
@@ 2 Preliminaries
$\mathcal{A}\,\mathcal{C}$ ::=:= $\{AC\,|\,A\in\mathcal{A},C\in\mathcal{C}\}$
$\mathcal{A}^{*}$ ::=:= $\{A^{*}\,|\,A\in\mathcal{A}\}\ \quad\text{with}\ *\in\{\top,-1\}.$
We let $\operatorname{co}(\mathcal{A})$ be the convex envelope of $\mathcal{A}$, and define $\lVert\mathcal{A}\rVert_{\sup}=\sup\{\lVert A\rVert\,|\,A\in\mathcal{A}\}$. It will be convenient to define for every $\mathcal{A}\subset\mathbb{R}^{n\times(p_{1}+p_{2})}$ the map acting between sets of matrices, which we still denote by $\mathcal{A}$, such that for every $\mathcal{X}\subset\mathbb{R}^{p_{1}\times p_{2}}$
$\mathcal{A}(\mathcal{X}){:=}\mathcal{A}\,\begin{bmatrix}\mathcal{X}\\ I_{p_{2}}\end{bmatrix}{:=}\big{\{}A_{1}X+A_{2}\,\big{|}\,[A_{1}A_{2}]\in\mathcal{A},X\in\mathcal{X}\big{\}},$ (3)
where $I_{p_{2}}$ is the identity matrix of dimensions $p_{2}\times p_{2}$.
For any integer $r\geq 1$ we set $[r]=\{1,\dots,r\}$. If $F\colon\mathbb{R}^{p_{1}+p_{2}}\to\mathbb{R}^{d}$ is differentiable, we denote by $F^{\prime}(x)\in\mathbb{R}^{d\times(p_{1}+p_{2})}$ the derivative of $F$ (its Jacobian) at $x$ and by $\partial_{1}F(x)\in\mathbb{R}^{d\times p_{1}}$ and $\partial_{2}F(x)\in\mathbb{R}^{d\times p_{2}}$ the partial derivatives of $F$ with respect to the first and second block of variables respectively. For a random vector $\xi\in\mathbb{R}^{d}$, we denote with $\mathbb{E}[\xi]$ its expectation and with $\operatorname{Var}[\xi]=\mathbb{E}\lVert\xi-\mathbb{E}[\xi]\rVert^{2}$ its variance. In our assumptions we will consider the class of so called definable functions, which includes the large majority of functions used for machine learning applications (see Appendix A).
@@ 2.1 Conservative Derivatives
Let $U\subset\mathbb{R}^{p}$ be an open set and $F\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ be a locally Lipschitz continuous mapping. We say that a set-valued mapping $D_{F}\colon U\rightrightarrows\mathbb{R}^{d\times p}$ is a conservative derivative (CD) of $F$, if $D_{F}$ has closed graph, nonempty compact values, and for every absolutely continuous curve $\gamma\colon[0,1]\to U\subset\mathbb{R}^{p}$ we have that, for almost every $t\in[0,1]$
@@
The function $F$ is called path differentiable if it admits a conservative derivative.
Conservative derivatives are extensively analyzed in (Bolte & Pauwels, 2021). Some key properties are that: (1) they are almost everywhere single-valued and equal to classical derivatives; (2) for path differentiable functions, the Clarke subgradient is the minimal conservative derivative up to a convex envelope; (3) chain rule holds for conservative derivatives; (4) locally Lipschitz definable mappings admit conservative derivatives. We also point out that – as it is usual for generalized derivatives – conservative derivatives are unique only up to a set of Lebesgue measure zero. This accounts for the fact that there are multiple ways to express a path differentiable function as a composition of others but applying the chain rule produces always valid CDs.
@@
Note that $\lVert\mathcal{A}\rVert_{\sup}=\operatorname{\it{e}}(\mathcal{A},\{0\})$. The excess satisfies several properties similar to the ones of a distance, even though it is not symmetric (see Lemma B.1). Similarly to (Scholtes, 2012) we give the following concept of piecewise continuity and smoothness (which is slightly more general than that given in (Bolte & Pauwels, 2021)).
@@
Let $F_{1},\dots,F_{r}\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ be continuous mappings defined on a nonempty open set $U$. A continuous selection of $F_{1},\dots,F_{r}$ is a continuous mapping $F\colon U\to\mathbb{R}^{d}$ such that for every $x\in U\colon F(x)\in\{F_{1}(x),\dots,F_{r}(x)\}$. In such case the active index set mapping is the set-valued mapping $I_{F}\colon U\rightrightarrows[r]$, with $I_{F}(x)=\{i\in[r]\,|\,F_{i}(x)=F(x)\}$. Moreover, if the $F_{i}$’s are differentiable we set $D_{F}^{s}\colon U\subset\mathbb{R}^{p}\rightrightarrows\mathbb{R}^{d\times p}$ such that
@@
Let $F\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ be a continuous selection of definable and continuously differentiable mappings $F_{1},\dots,F_{r}\colon U\to\mathbb{R}^{d}$. Then $F$ is definable if and only if $I_{F}\colon\mathbb{R}^{p}\rightrightarrows[r]$ is definable, and in such case $D_{F}^{s}$ is a conservative derivative of $F$.
We can also define partial conservative derivatives. If $p=p_{1}+p_{2}$ and $F\colon U\subset\mathbb{R}^{p_{1}+p_{2}}\to\mathbb{R}^{d}$, we have $D_{F}\colon U\rightrightarrows\mathbb{R}^{d\times(p_{1}+p_{2})}$ and we set $D_{F,1}\colon U\rightrightarrows\mathbb{R}^{d\times p_{1}}$ and $D_{F,2}\colon U\rightrightarrows\mathbb{R}^{d\times p_{2}}$ such that for $j\in\{1,2\}$
$D_{F,j}(x)=\big{\{}A_{j}\,|\,[A_{1},A_{2}]\in D_{F}(x)\big{\}}.$
Finally, we denote by $F^{\prime}(x)$ an arbitrary element of $D_{F}(x)$ and by $\partial_{1}F(x)\in\mathbb{R}^{d\times p_{1}}$ and $\partial_{2}F(x)\in\mathbb{R}^{d\times p_{2}}$ the first and second block component of $F^{\prime}(x)$ respectively, which yield the classical (partial) derivatives if $F$ is differentiable. By building on (Bolte et al., 2022, Lemma 3), we prove the following result (the proof is in Appendix B).
@@
Let $F\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ be a continuous definable selection of the definable Lipschitz smooth mappings $F_{1},\dots,F_{r}\colon U\to\mathbb{R}^{d}$. Let $L_{i}$ be the Lipschitz constant of $F^{\prime}_{i}$ and set $L=\max_{1\leq i\leq r}L_{i}$. Then for every $x\in U$, there exist $R_{x}>0$ such that for every $x^{\prime}\in U$
$\operatorname{\it{e}}(D_{F}^{s}(x^{\prime}),D_{F}^{s}(x))\leq L_{x}(x^{\prime})\lVert x-x^{\prime}\rVert,$ (6)
@@
$L_{x}(x^{\prime}):=\begin{cases}L&\text{if }\lVert x-x^{\prime}\rVert\leq R_{x}\\ L+M_{x}/R_{x}&\text{otherwise}\end{cases}\quad\text{and}\quad M_{x}:=\max_{i\in[m]}\min_{j\in I_{F}(x)}\lVert F^{\prime}_{i}(x)-F^{\prime}_{j}(x)\rVert.$
Note that in the smooth case ($r=1$), (6) corresponds to global $L$-smoothness (since $M_{x}=0$), while in general it is weaker. In particular, the quantity $L+\frac{M_{x}}{R_{x}}$ is well defined even when $F$ is not differentiable at $x$, but blows up when $x$ approaches a point of non-differentiability, e.g., for $\operatorname{ReLU}(x)=\max(0,x)$, $\lim_{x\to 0^{+}}M_{x}/R_{x}=\infty$, since if $x\neq 0$ $M_{x}=1$ and $R_{x}=\lvert x\rvert$, while for $x=0$, $M_{x}/R_{x}=0$ since $M_{x}=0$ and $R_{x}>0$ can be chosen arbitrarily.